OpenStudy (anonymous):

Which value of c makes the expression x^2-5x+c a perfect square trinomial? a)-5/2 b)5/2 c)25/4 d)25

hartnn (hartnn):

divide the co-efficient of x by 2 what u get ?

OpenStudy (anonymous):

-5/2 ?

hartnn (hartnn):

yup, now square it, what u get ?

OpenStudy (anonymous):

i got -1.25

hartnn (hartnn):

square of a number can never be negative...

hartnn (hartnn):

(-5/2)^2 = 25/4

OpenStudy (anonymous):

oh I see, thank u very much

hartnn (hartnn):

so, c= 25/4 ok, welcome ^_^